OpenStudy (loser66):

Let \(F:= \left[\begin{matrix}1\\1\\1\\1\\1\end{matrix}\right] +p\left[\begin{matrix}0\\0\\1\\1\\1\end{matrix}\right]+q\left[\begin{matrix}0\\1\\1\\1\\1\end{matrix}\right]+r\left[\begin{matrix}1\\1\\1\\1\\1\end{matrix}\right]\) Find a matrix A and vector x, b such that \(F:={x \in R^? |Ax=b}\) Please, help

OpenStudy (loser66):

Dimension of F is 3 , but \(\vec x \in \mathbb R^5\) How to go back to original system? I know how to go forward but don't know how to go backward.

OpenStudy (loser66):

I meant if I have a system of equation, after doing row reduce form, I can get this parametric system. Now, the exercise gives me the parametric system, I have to go back ward to find the system. :(

OpenStudy (mathmate):

.

OpenStudy (loser66):

Let \(\vec x=\left[\begin{matrix}x_1\\x_2\\x_3\\x_4\\x_5\end{matrix}\right] \)

OpenStudy (loser66):

then, \(x_1=1+r\\x_2= 1+q+r\\x_3=x_4=x_5=1+p+q+r\)

OpenStudy (loser66):

then what?

OpenStudy (mathmate):

It seems to me the way F is defined, the last three variables are not independent. so something like: 1 2 5 5 5 2 7 1 1 1 3 4 2 2 2 5 4 3 3 3 7 1 4 4 4 gives the REF as 1 2 5 5 5 0 1 -3 -3 -3 0 0 1 1 1 0 0 0 0 0 0 0 0 0 0 and gives the solution as shown.

OpenStudy (mathmate):

For b, you can try [13,12,13,18,20] (sum of all the coefficients)

OpenStudy (loser66):

How can you get it? the last three p, q , r vectors are independent. they span F, F is defined as \(F:= v_0 + pV_1+q V_2+rV_3\)

ganeshie8 (ganeshie8):

I think 'b' is the zero vector because the given particular solution is part of the null solution.

ganeshie8 (ganeshie8):

3 free variables in the general solution implies the rank is 5-3 = 2. That means the number of independent columns is also 2

ganeshie8 (ganeshie8):

Try this : A: 0 0 1 0 -1 0 0 0 1 -1 b: 0 0

OpenStudy (mathmate):

Notice that by arranging linear combinations, F can be made to read \(F:= \left[\begin{matrix}1\\1\\1\\1\\1\end{matrix}\right] +u\left[\begin{matrix}0\\0\\1\\1\\1\end{matrix}\right]+v\left[\begin{matrix}0\\1\\0\\0\\0\end{matrix}\right]+w\left[\begin{matrix}1\\0\\0\\0\\0\end{matrix}\right]\) From where I had the idea that the last three variable share the same value.

OpenStudy (loser66):

@ganeshie8 I would like to know how you get that matrix, please

ganeshie8 (ganeshie8):

Does it work ?

OpenStudy (loser66):

I don't see the link between F and the matrix you gave me :(

ganeshie8 (ganeshie8):

Let's do this on a quick teamviewer meeting if you still have it

OpenStudy (loser66):

You take control, right?

ganeshie8 (ganeshie8):

No.. You may just start a meeting and send me the id...

OpenStudy (loser66):

My ID 927 189 615

ganeshie8 (ganeshie8):

That's for controlling See if you can start a meeting

OpenStudy (loser66):

It says nothing :(

ganeshie8 (ganeshie8):

Meeting option will be there on right hand side of the teamviewer main window

OpenStudy (loser66):

From the right, it asks for partner ID

ganeshie8 (ganeshie8):

Right top side, I guess

OpenStudy (loser66):

This is what I have and it doesn't allow me to do anything

OpenStudy (loser66):

@ganeshie8 we work here, please.

ganeshie8 (ganeshie8):

I'm on mobile Can't type much sry

ganeshie8 (ganeshie8):

I can't open that docx file

OpenStudy (loser66):

That's ok then. I show you my work, you check , ok?

ganeshie8 (ganeshie8):

Sure

ganeshie8 (ganeshie8):

May I know why you think the column vectors must have 5 components ?

OpenStudy (loser66):

because the way they defined F, any vector in F has 5 components

OpenStudy (loser66):

\(F:= \left[\begin{matrix}1\\1\\1\\1\\1\end{matrix}\right] +p\left[\begin{matrix}0\\0\\1\\1\\1\end{matrix}\right]+q\left[\begin{matrix}0\\1\\1\\1\\1\end{matrix}\right]+r\left[\begin{matrix}1\\1\\1\\1\\1\end{matrix}\right]\)

ganeshie8 (ganeshie8):

F is not a matrix

ganeshie8 (ganeshie8):

F is the set of vectors that are solutions to Ax = b

OpenStudy (loser66):

oh yeah.

ganeshie8 (ganeshie8):

Hey try starting a meeting if you can

OpenStudy (loser66):

Oh, I think I got your given matrix Since the last 3 equation are the same, you have \(x_3=x_5\\x_4=x_5\) that gives you the matrix above

OpenStudy (loser66):

:) ganeshie8 , every time I try to connect you on teamviewer, I am so embarrass when struggling with the process to access into it. It belittles me. I fell as if I am the most stupid people in the world hahaha....... everybody can use it but I. As a result, my face turns red now and I just want to give up no matter what it is.

ganeshie8 (ganeshie8):

Haha thats okay Have u looked at the link I attached

OpenStudy (loser66):

yes, I did

ganeshie8 (ganeshie8):

You just have to click on Meeting, then start meeting

OpenStudy (loser66):

Yes, I did also. It is ask for meeting ID

ganeshie8 (ganeshie8):

Youre on the right tab. On the left side, Click start instant maating

OpenStudy (loser66):

It is m, but when I try to click on "join meeting", it says "it is not a valid meeting-ID

ganeshie8 (ganeshie8):

You're trying to join a meeting. You should start the meeting instead

OpenStudy (loser66):

AAAAAAAAAAHHH. I tried any button and got nothing. hahahaha.. .... ganeshie8, please ....... as my friend, don't ask me to go there now. I promise. I will master it next week with ....another one, not you. hehehe..... then, I will contact you to show the resutl.

ganeshie8 (ganeshie8):

Try once Its not rocket science. Not for you

OpenStudy (loser66):

Yes, sir. but later, not now. gtg.